OpenStudy (anonymous):

solve the equation. check your answer. m/20 = 5/4

OpenStudy (anonymous):

multiply both sides by 20 to get m by itself on the left side.

OpenStudy (anonymous):

multiply both sides by 20

OpenStudy (anonymous):

the 20 and the 5 right

OpenStudy (anonymous):

what about the 5?

OpenStudy (anonymous):

what side am I multiplying

OpenStudy (anonymous):

multiply 20 on both sides

OpenStudy (anonymous):

with the 4 and 5 I am not understanding

OpenStudy (anonymous):

\[20*\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

\[\frac{m}{20} = \frac{5}{4}\] \[m = \frac{5*20}{4}=\frac{100}{4}\]

OpenStudy (anonymous):

i got 25

OpenStudy (anonymous):

20 *5/4=25